OpenStudy (megannicole51):

Throughout much of the 20th century, the yearly consumption of electricity in the US increased exponentially at a continuous rate of 7% per year. Assume this trend continues and that the electricity energy consumed in 1900 was 1.4 million megawatt-hours. (a) Write an expression for yearly electricity consumption as a function of time, t, in years since 1900. (b) Find the average yearly electrical consumption throughout the 20th century. (c) During what year was electrical consumption closest to the average for the century? (d) Without doing the calculator for part (c), how could you have

OpenStudy (megannicole51):

predicted which half of the century the answer would be?

OpenStudy (anonymous):

of something increases exponentially. you are going to be working with something^(t) where t is your time in years. 7% increase means that one year later it is 107% of the last year. Try and find an exponent involving t that will increase your base by 7% when t is 1 (i.e 1 year later)

OpenStudy (megannicole51):

im sorry i dont really understand can you elaborate

OpenStudy (anonymous):

as Taplin44 said, you need an equation:\ y(t) = Ae^(bt) where y(0) = A (energy at 1900 which is given)

OpenStudy (anonymous):

as for b, you want this factor e^b (t=1) to be 1.07 so that after 1 year you will have a grow of 107% 1.07A

OpenStudy (anonymous):

so now you can find b : e^b = 1.07

OpenStudy (anonymous):

sorry for my English.

OpenStudy (megannicole51):

do u mean the exponent b or the second question?

OpenStudy (anonymous):

Yes, he did.

OpenStudy (anonymous):

yes i did lol

OpenStudy (anonymous):

Coolsector is so cool!

OpenStudy (megannicole51):

so then how would i do (c)?

OpenStudy (anonymous):

you found the average ?

OpenStudy (anonymous):

Yeah, did you find the average?

OpenStudy (anonymous):

all you have to do is average = Ae^(bt) and now find t

OpenStudy (anonymous):

You can do this! I believe in you1

OpenStudy (megannicole51):

how do i bring down the exponent? its been a long night of calc and my brain is fried

OpenStudy (anonymous):

you mean how you find b using e^b = 1.07 ?

OpenStudy (megannicole51):

yes

OpenStudy (anonymous):

take ln from both sides

OpenStudy (anonymous):

Yes, Coolsector is correct, he is never wrong!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (megannicole51):

asdfasdf54514....get a life.

OpenStudy (megannicole51):

b=ln1.07

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

LOL... see, he doesn't lie!!!!!!

OpenStudy (megannicole51):

so thats the average?

OpenStudy (anonymous):

Block me please. :) Actually never mind, I will leave.How does one untag himself?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

now we are only working on finding the expression for (a)!

OpenStudy (anonymous):

im not talking about (b). my bad for calling the parameter b

OpenStudy (anonymous):

but it is only part of the expression

terenzreignz (terenzreignz):

@Coolsector a question, do we really have to go through the part where we have \[\Large Ae^{bt}\] with the constant e instead of just using \[\Large A(1.07)^t\]?

OpenStudy (anonymous):

no

OpenStudy (megannicole51):

im kinda confused

terenzreignz (terenzreignz):

Sorry @megannicole51 I was only asking Coolsector... don't worry, you guys seem to be on the right track :D

OpenStudy (megannicole51):

oh okay lol then keep going I'm super tired and this is my last problem! :)

OpenStudy (anonymous):

so ill try again. we are now looking for the expression they ask in (a): we want exponential expression of the form: y(t) = Ae^(Bt) y(t) will tell us the energy consumption at time t in years after 1900 so that y(0) = A will be the energy consumption at 1900!

OpenStudy (anonymous):

so can you tell me what A is ?

OpenStudy (megannicole51):

Y(0)?

OpenStudy (anonymous):

y(0) = the energy consumption at the year 1900

OpenStudy (megannicole51):

im not sure what A is then

OpenStudy (anonymous):

y(0) = A = the energy consumption at the year 1900 it is given in the question

OpenStudy (megannicole51):

so i was right?

OpenStudy (anonymous):

where ?

OpenStudy (megannicole51):

u asked what A was

OpenStudy (anonymous):

so what is it ?

OpenStudy (megannicole51):

idk.

OpenStudy (anonymous):

again, the expression we are working on y(t) = Ae^(bt) y - the energy consumption since 1900 (when t=0 it tells us the energy consumption at 1900 when t =1 it tells us the energy consumption at 1901 etc)

OpenStudy (anonymous):

so y(0) = A and it is given in the question : 1.4 million megawatt-hours

OpenStudy (anonymous):

so now we have: y(t) = 1.4e^(bt) and it satisfies our first condition : y(0) = 1.4

OpenStudy (anonymous):

(y in units of million megawatt-hours)

OpenStudy (anonymous):

understand so far ?

OpenStudy (kropot72):

The equation for (a) is: \[Yearly\ consumption=1.4e ^{0.07t}\] (b) \[Average\ thru \ 20th\ century=\frac{\int\limits_{t=0}^{t=100}1.4e ^{0.07t}.dt}{100}\]

OpenStudy (megannicole51):

okay go on

OpenStudy (anonymous):

well i dont understand what is the point in giving the answers but whatever

OpenStudy (kropot72):

Notice that using the equation for (a) the growth after 1 year is 7.251% not 7%, the reason being that the growth is stated to be exponential and continuous. The formula is derived using calculus. Can you evaluate the definite integral in the equation for (b)?

OpenStudy (kropot72):

@Coolsector You posted "..........to be 1.07 so that after 1 year you will have a grow of 107%". Note that after 1 year the consumption has grown to 107.251% of the initial consumption. I intervened to try to clarify matters.

OpenStudy (anonymous):

if we could keep all the numbers of b from e^b = 1.07 we would get a growth of 7%

OpenStudy (anonymous):

the values of y for t = 1,2,3,4.. would be accurate if we could.

OpenStudy (kropot72):

@Coolsector You are confusing the "continuous growth rate of 7% per year" with the annual growth and expecting the consumption after 1 year to be 107% of the initial consumption. However this expectation is not correct. You will find a good explanation of exponential growth here. http://math.ucsd.edu/~wgarner/math4c/textbook/chapter4/expgrowthdecay.htm

OpenStudy (anonymous):

i understand what you are trying to say. but still every year the consumption grows at 7% for t not integer we will get different numbers but when t is integer we will get every time growth of 7% since it is like "the end of the year" and i think there is nothing to do with that 107.251% that you talked about. this is because of a different reason - the reason that you rounded up b to 0.07.

OpenStudy (anonymous):

anyway this is not really matters here.

OpenStudy (kropot72):

You are confusing 'Discrete exponential growth' and 'Continuous exponential growth'. The question deals with continuous exponential growth. The difference certainly does matter here. The derivation of the formula for continuous exponential growth is given here. http://mathworld.wolfram.com/ExponentialGrowth.html

OpenStudy (anonymous):

im not confusing anything. the way i was solving it is correct (this is what you did in fact)

OpenStudy (anonymous):

the expression we got is continuous but it still give the right values for every year

OpenStudy (kropot72):

What expression did you get?

OpenStudy (anonymous):

i wrote up there "you need an equation:\ y(t) = Ae^(bt) where y(0) = A (energy at 1900 which is given) as for b, you want this factor e^b (t=1) to be 1.07 so that after 1 year you will have a grow of 107% 1.07A so now you can find b : e^b = 1.07"

OpenStudy (anonymous):

continuous expression can be still correct for specific discrete values. such as here.

OpenStudy (kropot72):

But your value for b is incorrect. The growth rate is 0.07 per year, so b must equal 0.07. You have contrived a value for b to make the consumption after 1 year be 107% of the initial consumption. That is incorrect and you are not following the theory in the link that I posted.

OpenStudy (anonymous):

my value for b is 0.067658648.. which gives e^b = 1.07 which is the exact value

OpenStudy (anonymous):

because you rounded b into 0.07 you got a higher precent

OpenStudy (anonymous):

there is nothing we can do about it - we have to round. but this is the reason.

OpenStudy (kropot72):

No, I did not round the value. I used the decimal equivalent of 7% as given in the question.

OpenStudy (anonymous):

ok

OpenStudy (kropot72):

As explained in the link, continuous compounding as used in finance is basically the same. In the case of annual compounding at an annual interest rate of 7% a $100 initial investment is worth $107 after the first year. However with continuous compounding at an annual interest rate of 7%, the amount after 1 year is \[A=100e ^{0.07}=$107.251\]

OpenStudy (anonymous):

well i still dont understand why you think determining 0.07 is better than finding this 0.067.. if i look at the examples at the link you gave they do what i did

OpenStudy (anonymous):

but really i dont care for it. rounding this number would give me 0.07 so i would go with it anyway

OpenStudy (anonymous):

anyway. look at higher numbers like if the growth was 60% what would you do then ? say that the factor is 0.6?! this is completely wrong

OpenStudy (anonymous):

for 60% growth the factor should be about 0.47

OpenStudy (anonymous):

please answer for my last two comments as well.

OpenStudy (kropot72):

Your reasoning is incorrect. You need to understand the ordinary differential equation that leads to the formula for exponential growth. You are taking a figure for growth and trying to fit it to discrete values of time (years). Taking 60% growth in consumption we get: \[1.4\times1.6=1.4e ^{0.07t}\] Solving for t gives t = 6.714 years Note that fractions of a year are valid for continuous exponential growth.

OpenStudy (anonymous):

no i said instead of 7% growth 60% growth in one year. find the factor b then.

OpenStudy (anonymous):

according to you it should be 0.6 which cant be right

OpenStudy (kropot72):

The growth factor remains constant when applying the equation for (a) in the question: \[Yearly\ consumption=1.4e ^{0.07t}\]

OpenStudy (kropot72):

You posted "according to you it should be 0.6" I made no such statement.

OpenStudy (anonymous):

im talking about different question now . im asking you, what would be the growth factor if the the rate was 60% instead of 7%

OpenStudy (anonymous):

according to your logic it should be 0.6 this cant be right

OpenStudy (kropot72):

Are you talking about a situation where continuous annual growth is 60%. If so, assuming an initial quantity of 100 units, the quantity after 1 year would be \[N=100e ^{0.6}=182.2\ units\]

OpenStudy (anonymous):

so i say, the factor should be 0.47 e^(0.47)

OpenStudy (anonymous):

N=100e^0.47 = 159.99 this is much better

OpenStudy (kropot72):

Your error is in expecting the amount after 1 year to be 160 units. You have not grasped the difference between discrete exponential growth and continuous exponential growth.

OpenStudy (anonymous):

even in your links there is nothing that shows that your way is better

OpenStudy (kropot72):

Give me details of where you found "if i look at the examples at the link you gave they do what i did"

OpenStudy (anonymous):

all the examples at the first link they find the factor using my method

OpenStudy (kropot72):

I do not agree that any of the examples follow your method. Your method is based on a hybrid of discrete and continuous growth formulas. Look at the explanation and examples here: http://people.stern.nyu.edu/wsilber/Continuous%20Compounding.pdf

OpenStudy (anonymous):

ithink, that this approximation (which is the same as rounding) only good for low numbers

OpenStudy (kropot72):

Note the following quote from the last link: " In our context, this means that if $1 is invested at 100% interest, continuously compounded, for one year, it produces $2.71828 at the end of the year. " Accord to your reasoning the amount produced should be exactly $2.00.